An oxide of antimony is found to contain 24.73% oxygen. What is the empirical formula?


1
Expert's answer
2019-09-19T05:11:14-0400

"\frac{24.73}{M_O}:\frac{100-24.73}{M_{Sb}}=\frac{24.73}{16.00}:\frac{75.27}{12.176}=1.55:0.62=5:2"

Sb2O5